OpenStudy (anonymous):

Find the area of a sector with the given central angle Theta in radius r: Theta= 2(in radians) and r = 10 inches Very confused how they got 100 sq inches

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

We know:\[A _{sector}=(1/2)r^2\theta\] Keep in mind that theta in this case must be in radians, not degrees. So,\[A _{sector}=(1/2)10^2(2)\]\[A _{sector}=(1/2)200\] \[A _{sector}=100in^2\]

OpenStudy (anonymous):

I can't find that formula in my book hmm... and it says it falls under the category of arc length, linear and angular velocity chapter. (This is a trig class). any way of doing that with s=r*theta or v= s/t or w= theta/t? or is it far more basic than that?

OpenStudy (anonymous):

What is s, v, and w in those formulas?

OpenStudy (anonymous):

Also "t"

OpenStudy (anonymous):

arc length linear velocity and angular velocity

OpenStudy (anonymous):

t is just a length on the unit circle

OpenStudy (anonymous):

can you just explain to me the forumla a little like where you got 1/2 what r^2 represent etc

OpenStudy (anonymous):

hm, I've never seen sector area calculated via kinematics. Deriving a formula for sector area from those 3 equations seems very difficult since none of them have an area variable in them. I remember the basic formula for the area of a sector is\[A _{sector}=(1/2)r^2\theta\]Where r is the radius and theta is the angle of the sector.

OpenStudy (anonymous):

No need for all this complication. The area of the circle is pi r^2 ie a 100 pi Your piece of it is 2/2pi = 1/pi radians (there are 2pi radians in the whole circle) SO 100 pi + 1/pi = 100 That's it.

OpenStudy (anonymous):

100 pi * 1/pi = 100 (not +)